TheSmartOne (thesmartone):

10000 = 4000(1.04)^t Solve for t.

OpenStudy (anonymous):

\[\begin{align*} 10000&=4000(1.04)^t\\\\ \frac{10000}{4000}&=\frac{4000}{4000}(1.04)^t\\\\ \frac{10}{4}&=(1.04)^t\\\\ \log_{1.04}\frac{10}{4}&=\log_{1.04}(1.04)^t\\\\ \log_{1.04}\frac{10}{4}&=t\log_{1.04}1.04\\\\ \log_{1.04}\frac{10}{4}&=t \end{align*}\]

OpenStudy (anonymous):

Use the change of base formula: \[\large \log_ba=\frac{\ln a}{\ln b}\]

OpenStudy (anonymous):

To be fair, I already have shown you how to get the answer. It's a matter of using a calculator now. http://www.wolframalpha.com/input/?i=ln%2810%2F4%29%2Fln%281.04%29

OpenStudy (jdoe0001):

\(\Large \bf { {\color{blue}{ b}}={\color{brown}{ a}}^y\implies log_{\color{brown}{ a}}{\color{blue}{ b}}=y\qquad thus \\ \quad \\ {\color{blue}{ 10000}} = {\color{brown}{4000(1.04)}}^t\implies ? }\)

OpenStudy (horsegirl27):

;)